# Auxiliary task demands mask the capabilities of smaller language models

### v1

@@ Abstract
Developmental psychologists have argued about when cognitive capacities such as language understanding or theory of mind emerge. These debates often hinge on the concept of “task demands” – the auxiliary challenges associated with performing a particular evaluation – that may mask the child’s underlying ability. The same issues arise when measuring the capacities of language models (LMs): performance on a task is a function of the model’s underlying competence, combined with the model’s ability to interpret and perform the task given its available resources. Here, we show that for analogical reasoning, reflective reasoning, word prediction, and grammaticality judgments, evaluation methods with greater task demands yield lower performance than evaluations with reduced demands. This “demand gap” is most pronounced for models with fewer parameters and less training data. Our results illustrate that LM performance should not be interpreted as a direct indication of intelligence (or lack thereof), but as a reflection of capacities seen through the lens of researchers’ design choices.
@@ 2 Background and Related Work
An evaluation with fewer task demands is not necessarily easier or more lenient. Instead, a better gloss might be that it is a purer task, in the sense that it better measures the targeted construct. Consider an evaluation of arithmetic ability: the size of the numbers being manipulated and the range of operations would likely change the overall difficulty of the evaluation. In contrast, the use of word problems, pictorial representations, or framing in terms of real-world objects would all change the task demands of the evaluation, even if the difficulty of arithmetic components stayed constant. Some of these changes might serve to make the task easier – for example, if they engaged familiar content (Dasgupta et al., 2022) – while others might make the task harder (at least for some models).
@@ 3 Methods
Our main research question is whether less capable language models are more sensitive to task demands. If this is the case, then we would expect to find the pattern in Figure 1C: the performance gap between low- and high-demand tasks should be smaller for more capable models. To avoid conflating the targeted cognitive construct with a specific contrast of evaluation methods, we perform each evaluation comparison on two different domains. In the remainder of this section, we discuss the evaluation contrasts, cognitive constructs, and models tested in our experiments (see Table 1 for summary).
@@ 3.1 Evaluation contrasts
Production (high demands) vs. Forced choice (low demands) Analogical reasoning Webb et al. (2023) Predict next digit(s) in sequence [5 9 3] [8 9 2] [1 9 7] $\backslash\n$ [8 4 7] [1 4 3] [5 4 2] $\backslash\n$ [1 2 2] [5 2 7] [
@@
We focus on two contrasts of high- and low-demand evaluation methods: (1) production (high) versus forced choice (low), and (2) metalinguistic judgments (high) versus direct probability measurements (low). Each of these four evaluation methods has previously been applied to make claims about cognitive abilities in language models.
@@
It is not obvious that this asymmetry would also apply for LMs, for which the task of generating linguistic output – which may impose auxiliary demands on children, including planning motor actions associated with speech production – is the primary rationale. But even without these demands, forced choice requires less precise estimates of a probability distribution than correct production. A correct answer might be reliably preferred over incorrect answers in probability space, while not being generated because of auxiliary factors such as the decoding mechanism, or simply having a low absolute probability. [^2] Forced choice evaluation does require researchers to design a set of answer options, which may introduce unintended artifacts (e.g., content or frequency effects). In line with this view, West et al. (2024) find that generative (production) performance is better than discriminative (forced choice) performance across a range of models and tasks. However, this finding could be due to the complexity of the answer options in their study, which were designed to include adversarial “hard negatives”. More broadly (and in keeping with our general theme), researcher goals (in this case, adversarial testing as opposed to capacity discovery) can modulate the difficulty of any evaluation method.
@@ 3.3 Models
This analysis is not meant to draw direct parallels between development in children and scaling in LMs: we don’t claim, for example, that Llama-2 7B is more like a young child and Llama-2 70B is more like an adult. Instead, we use size as a practical way of parameterizing the “capacity” of a model, which might affect its ability to overcome auxiliary task demands.
@@ 4 Results
Figure 2 shows the results for our first tested contrast: production vs. forced choice. We first briefly discuss raw task performance. For the analogical reasoning domain, larger models generally achieve higher performance (compared to smaller models in the same family) under both evaluation methods (Figure 2A). Interestingly, we do not find this pattern for the reflective reasoning domain (Figure 2C), because the larger models have a strong preference for intuitive responses, whereas the small models slightly prefer correct >>> intuitive but have a much stronger preference for atypical answers (see Hagendorff et al. 2023 for details).

### v2

@@ Abstract
Developmental psychologists have argued about when cognitive capacities such as language understanding or theory of mind emerge. These debates often hinge on the concept of “task demands” – the auxiliary challenges associated with performing a particular evaluation – that may mask the child’s underlying ability. The same issues arise when measuring the capacities of language models (LMs): performance on a task is a function of the model’s underlying knowledge, combined with the model’s ability to interpret and perform the task given its available resources. Here, we show that for analogical reasoning, reflective reasoning, word prediction, and grammaticality judgments, evaluation methods with greater task demands yield lower performance than evaluations with reduced demands. This “demand gap” is most pronounced for models with fewer parameters and less training data. Our results illustrate that LM performance should not be interpreted as a direct indication of intelligence (or lack thereof), but as a reflection of capacities seen through the lens of researchers’ design choices.
@@ 2 Background and related work
While the psychology literature does not provide precise definitions of what counts as a task demand, informally a task demand can be thought of as a feature of the task that imposes challenges unrelated to the capacity that the task is designed to measure. For humans (and particularly in children), these challenges might be related to processes such as memory, attention, or inhibition, which are involved in performing the task but separate from the target capacity. It is unclear how these kinds of challenges manifest in LM computation, and our work aims to explore several such possibilities.
An evaluation with fewer task demands is not necessarily easier or more lenient. Instead, a better gloss might be that it is a purer task, in the sense that it better measures the targeted construct. Consider an evaluation of arithmetic ability: the size of the numbers being manipulated and the range of operations would likely change the overall difficulty of the evaluation. In contrast, the use of word problems, pictorial representations, or framing in terms of real-world objects would all change the task demands of the evaluation, even if the difficulty of arithmetic components stayed constant. Some of these changes might serve to make the task easier – for example, if they engaged familiar content (Lampinen et al., 2024) – while others might make the task harder (at least for some models).
@@ 3 Methods
Our main research question is whether less capable language models are more sensitive to task demands. If this is the case, then we would expect to find the pattern in Figure 1C: the performance gap between low- and high-demand tasks should be smaller for more capable models. In the remainder of this section, we discuss the evaluation contrasts, cognitive constructs, and models tested in our experiments (see Table 1 for summary).
@@ 3.1 Evaluation contrasts
[…]
@@
We focus on two evaluation contrasts that are relevant to most behavioral LM evaluations. When designing an evaluation, researchers have the choice to assess the LM in a generative (production) or discriminative (forced choice) setting. Independently, there is also the choice of formulating the task as a prompt or a direct measurement of string probabilities. Both of these decision points can be formulated as a contrast between methods with higher and lower levels of task demands, which we discuss in more detail below.
@@
It is not obvious that this asymmetry would also apply for LMs, for which the task of generating linguistic output – which may impose auxiliary demands on children, including planning motor actions associated with speech production – is the primary rationale. But even without these demands, forced choice requires less precise estimates of a probability distribution than correct production. Conditioned on a context (e.g., the statement of a reasoning problem), the LM’s ability to solve the task is encoded in its distribution over subsequent tokens. This distribution is the same, regardless of whether the LM is evaluated through forced choice or production. Production, which involves sampling an output from the distribution, might not demonstrate the model’s abilities because of auxiliary factors (e.g., the decoding method or low absolute probabilities), whereas targeted comparisons of answer options might reveal distinctions that are present in the model’s probability distribution but would not be sampled. In this sense, production and forced choice can target the same cognitive construct, while imposing different kinds of auxiliary demands. [^2] Forced choice evaluation does require researchers to design a set of answer options, which may introduce unintended artifacts (e.g., content or frequency effects). In line with this view, West et al. (2024) find that generative (production) performance is better than discriminative (forced choice) performance across a range of models and tasks. However, this finding could be due to the complexity of the answer options in their study, which were designed to include adversarial “hard negatives”. More broadly (and in keeping with our general theme), researcher goals (in this case, adversarial testing as opposed to capacity discovery) can modulate the difficulty of any evaluation method.
@@ 3.2 Cognitive domains
For the production versus forced choice contrast, we evaluated LMs in two cognitive domains: analogical reasoning and reflective reasoning. Reasoning abilities have been of great recent interest in LM evaluation (e.g., Kojima et al., 2022; Webb et al., 2023; Chang & Bergen, 2024), and the test sets we used feature well-motivated answer option sets that enable forced choice evaluation.
@@
For the metalinguistic judgment versus probability contrast, we evaluated LMs on word prediction and grammaticality judgment because they are the clearest cases where metalinguistic prompting differs from established ways of using direct measurements.
@@
Note that we only tested a single prompt for each of the metalinguistic evaluation tasks. For results on a broader set of prompts, we refer readers to Hu & Levy (2023). They tested models on 3 different metalinguistic prompts in both English and Chinese, and found similar qualitative findings across prompt types and languages.
@@ 3.3 Models
This analysis is not meant to draw direct parallels between development in children and scaling in LMs: we don’t claim, for example, that Llama-2 7B is a model of a young child and Llama-2 70B is a model of an adult. Instead, we use size and training as practical proxies for how LMs might gain resources that help with overcoming task demands.
@@ 4 Results
Figure 2 shows the results for our first tested contrast: production vs. forced choice. We first briefly discuss raw task performance. For the analogical reasoning domain, larger models generally achieve higher performance (compared to smaller models in the same family) under both evaluation methods (Figure 2A). Interestingly, we do not find this pattern for the reflective reasoning domain (Figure 2C), because the larger models have a strong preference for intuitive responses, whereas the small models slightly prefer correct $>$ intuitive but have a much stronger preference for atypical answers (see Hagendorff et al. 2023 for details).
@@
To formally test our demand gap prediction (visualized in Figure 1C) – i.e., that smaller models are differentially affected by task demands – we analyzed the interaction between model size and task demands. We used the lme4 package in R to fit the following mixed effects model for each of the domains with binary responses (analogical reasoning, reflective reasoning, and grammaticality judgment):
[…]
We fit the following model for the word prediction domain:
[…]
We found a significant interaction between model size and evaluation method for each domain except reflective reasoning (analogical reasoning: $z=2.631,p=0.009$; grammaticality judgment: $z=1.995,p=0.0461$; word prediction: $t=7.654;p=0.005$).
@@
Again, we fit generalized linear models testing for an interaction between training steps and evaluation method. Here, there is no model family grouping factor. We fit the following models for analogical reasoning and word prediction, respectively:
[…]
For both domains, we found a significant interaction between the number of training steps and evaluation method (analogical reasoning: $z=2.639,p=0.008$; word prediction: $t=24.41;p<0.0001$).